OpenStudy (deleonbrrj):

to prove that 5^3/2 = 10√5 ???

OpenStudy (anonymous):

That isnt true. \[5^{\frac{3}{2}}=\left(5^{\frac{1}{2}}\right)^3=5^{\frac{1}{2}}\cdot 5^{\frac{1}{2}}\cdot 5^{\frac{1}{2}}\]\[=5^{\frac{1}{2}+\frac{1}{2}}\cdot 5^{\frac{1}{2}}=5^1\cdot 5^{\frac{1}{2}}=5\sqrt{5}\]

OpenStudy (deleonbrrj):

perfect!!!!! thanks!!!